OpenStudy (anonymous):

what is x^5/x^2 simplified

OpenStudy (texaschic101):

exponent rule : x^m / x^n = x^(m-n) so.. x^5 / x^2 = x^(5-2) = x^3

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

um wait that's not 1 of the answers the answer are A 1,048,578 B 16,384 C 64 and D 16

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

it can only be x^3 unless you know what x is, we cannot go any further

OpenStudy (anonymous):

wait wait wait sry u were right I have the wrong question

OpenStudy (anonymous):

here what is 4^5/4^2 simplified that's the question

OpenStudy (texaschic101):

4^5/4^2 = 4^(5-2) = 4^3 = 64

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

lol...no problem :)